Editorial Board

Editor-in-Chief: Terrence Sejnowski, The Salk Institute

and University of California at San Diego

Advisory Board:

Shun-ichi Amari, University of Tokyo,  Teuvo Kohonen, University of Helsinki,
Japan  Finland
Michael Arbib, University of Southern Carver Mead, California Institute of
California  Technology
Jean-Pierre Changeux, Institut Pasteur, Tomaso Poggio, Massachusetts Institute of
France  Technology
Leon Cooper, Brown University  Wilfrid Rall, National Institutes of Health
Jack Cowan, University of Chicago  Werner Reichardt, Max-Planck-Institut für
Jerome Feldman, International  Biologische Kybernetik
Computer Science Institute  David A. Robinson, Johns Hopkins
Stephen Grossberg, Boston University  University
John Hopfield, California Institute of  David Rumelhart, Stanford University
Technology  Bernard Widrow, Stanford University

Action Editors:

Joshua Alspector, Bell Communications  Richard Lippmann, MIT Lincoln Laboratories
Research  Stephen Lisberger, University of Califor-
Richard Andersen, Massachusetts  nia at San Francisco
Institute of Technology  James McClelland, Carnegie-Mellon Uni-
James Anderson, Brown University  versity
Dana Ballard, University of Rochester  Graeme Mitchison, Cambridge University,
Harry Barrow, University of Sussex  England
Andrew Barto, University of Massachu-  David Mumford, Harvard University
setts  Erkki Oja, Lappeenranta University, Finland
Rodney Brooks, Massachusetts  Andras Pellionisz, New York University
Institute of Technology  Fernando Pineda, Jet Propulsion Laboratory,
Gail Carpenter, Northeastern University  Pasadena, California
Patricia Churchland, University of  Demetri Psaltis, California Institute of
California at San Diego  Technology
Gary Dell, University of Rochester  Idan Segev, Hebrew University, Israel
Jeffrey Elman, University of California  Allen Selverston, University of California
at San Diego  at San Diego
Nabil Farhat, University of Pennsylva-  Gordon Shepherd, Yale University
nia  Vincent Torre, Universita di Genova, Italy
Francois Fogelman-Soulie, Paris, France  David Touretzky, Carnegie-Mellon Uni-
Peter Getting, University of Iowa  versity
Ellen Hildreth, Massachusetts Institute  Roger Traub, IBM Yorktown Heights
of Technology  Les Valiant, Harvard University
Geoffrey Hinton, University of Toronto,  Christoph von der Malsburg, University
Canada  of Southern California
Bernardo Huberman, Xerox, Palo Alto  David Willshaw, Edinburgh University,
Lawrence Jackel, AT&T Bell Laboratories  Scotland
Scott Kirkpatrick, IBM, Yorktown Heights John Wyatt, Massachusetts Institute of
Christof Koch, California Institute of  Technology
Technology  Steven Zucker, McGill University, Canada